MODS Software Issue Trak Status

IT Detail Status
7018 setxy returns bad error number  
6939 synchronous offset timeout problem binocular issue - solved with sequencer?
6645 guide star on MODS1-Agw fainter than MODS2 at the moment, I don't think we have any ideas here
6570 timeouts, communications glitches or signs of race conditions mselect busy or already in progress messages
6543 MODS1 Red IMCS lock timing out problem with zero points?
not a problem like 5767, with the laser not coming on
6502 MODS2R acq images show mask motion transient issue, closed
6501 modsAlign XPA$ERROR invalid command name  
6500 MODS1R lagging RP closed, but has been reopened in Nov-2017
5931 MODS should always send guide star magnitude to GCS enhancement
5861 MODS1 Red exposure control hung on readout  
5713 MODS Exposure Time Not Updating Between ACQ and OBS scripts Olga thinks operator error on abort
5395 MODS AGw XY stages reporting incorrect coordinates ? This is still a problem; see also 6541
5303 Suggested changes/reorganization of MODS image headers enhancement
RP deferred until LBTO FITS spec is updated.
5222 MODS1 Red and Blue exposure controls hung on "Writing image..." enhancement
Need the script engine to catch PI_NAME > 68 characters
5175 MODS2 red IC error state this is very old and should be closed
5122 MODS1 AGw not where expected same as 5395 - reporting zero
5087 mods1data froze  
5086 IC program spontaneously quit on MODS1 Red computer M1.RC probably should be closed - not the computer issue anymore
4814 MODS red channel quadrant 1 readout malfunctioning should be closed, was fixed by swapping the sequencer board
4716 Vignetting in MODS1 Red Channel closed, nothing can be done about it
4586 MODS abort or retry causing subsequent script errors just needs clarification in the troubleshooting notes

What's the scoop with the bad readout ITs?
6386 MODS2 B bad readout  
6051 intermittent bad rows on MODS1  
5862 MODS1 Red bad readout  
5555 MODS2 Red CCD has a noisy upper right quadrant  
5544 mods1r image with 5 rows with constant value 32768 ADU  
4995 Noise in MODS1 Blue channel quadrant 2  
Topic revision: r8 - 06 Mar 2018, KelleeSummers
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback